Hello, may you please help me solve:
Is it more profitable to receive $1000 now or $1700 in nine years? Assume that money can earn 6% interest compounded annually.
Thank you!

Answer by nerdybill(7384)

Is it more profitable to receive $1000 now or $1700 in nine years? Assume that money can earn 6% interest compounded annually.
.
$1000 after 6% interest compounded annually for nine years:
A = P(1 + r)^t
A = 1000(1 + .06)^9
A = 1000(1.06)^9
A = 1000(2.172)
A = $2171.89
.
Answer: Take the $1000 now and invest it at 6% compounded annually